In some applications dealing with very high speeds, the velocity is measured
in mm/
1
Expert's answer
2021-08-30T01:54:53-0400

"a) \ 25mm/\mu{s}=\dfrac{25\times{10^{-3}}}{10^{-6}}=25000m/s"

"b)\ 
\dfrac{25000\times{0.000621}}{\frac{1}{60\times{60}}}=55890mi/h"